OpenStudy (anonymous):

The depth d(x), in feet, of a circular swimming pool at a point x feet from the edge is represented by the function below. The function d of x equals 1 + x/8 If the depth varies from 1 foot to 4 feet, which inequality describes the domain of the function? 0 ≤ x ≤ 8 0 ≤ x ≤ 24 0 ≤ x ≤ 40 1 ≤ x ≤ 8

OpenStudy (anonymous):

@FriedRice @tcarroll010 @shkrina @UnkleRhaukus @ivettef365

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

@mathslover Please help anyone ?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

d(x) = 1 to d(x) = 4 ....

OpenStudy (amistre64):

\[1\le d(x)\le 4\] \[1\le 1 + x/8 \le 4\] \[0\le x/8 \le 3\] \[0\le x \le 24\]

OpenStudy (anonymous):

ok thanks know i get it

OpenStudy (amistre64):

youre welcome ...